Please answer all please only
If you know it.
Directionsi Write out your legible answers in the space provided. 1. If a wright is hung from a string which has a teusion of 100 N, whuat i the mass of the weight 2. A rocket moving with a constant accleration that has a mass of 45 kg and an initial velocity of Om/s traveled 300 m in 10 . What is the force of the rocket? Chate the anavity of s of10 and a dianacter of 1171 10 kn G-6.673 x 10 N m2/kg
Answer #1

1) weight = 100 N

mass = weight / acceleration due to gravity

= 100/9.8 = 10.2 kg

2) S = ut + 0.5at^2

300 = 0 + 0.5*a*10^2

a = 6 m/s2

F = ma = 45*6 = 270 N

3) g = GM/r^2

= 6.67*10^-11*5.68*10^26 / (1.17*10^8/2)^2

= 11.07 m/s2